OpenStudy (pulsified333):

Two fair die are rolled. What is the probability of one of the dice showing 5 or the sum being at least 8?

OpenStudy (pulsified333):

well I know the sum of 8's probability is 5/36

OpenStudy (pulsified333):

right so wouldn't it be 1/6?

OpenStudy (anonymous):

+ the probability of at least 8

OpenStudy (pulsified333):

so would it be 6/36?

OpenStudy (pulsified333):

you mean 2/36?

OpenStudy (pulsified333):

so the answer would be 7/36?

OpenStudy (anonymous):

sorry can you forget about what I just said?

OpenStudy (pulsified333):

yeah

OpenStudy (pulsified333):

why?

OpenStudy (anonymous):

for at least one to be a 5 would be 11/36

OpenStudy (anonymous):

(for each dice)( 1/6 + 1/6) - 1/36 (because there's a double) = 11/36

OpenStudy (pulsified333):

so the answer is 11/36?

OpenStudy (anonymous):

+ at least 8

OpenStudy (pulsified333):

so 11/36+16/36

OpenStudy (anonymous):

- the mutual part

OpenStudy (pulsified333):

-1?

OpenStudy (pulsified333):

13/18?

OpenStudy (anonymous):

-6/36

OpenStudy (anonymous):

21/36

OpenStudy (pulsified333):

why -6/36

OpenStudy (anonymous):

because (5,1)(5,2)(5,3)(3,5)(2,5)(1,5) are included in both probabilities.

OpenStudy (pulsified333):

oh let me check

OpenStudy (anonymous):

I hope that's right.

OpenStudy (pulsified333):

would it be 7/12? simplified

OpenStudy (anonymous):

yes

OpenStudy (pulsified333):

I did it backwards the sum of being at least 8 would be 17/36

OpenStudy (pulsified333):

my bad i didn't see the at least 8

OpenStudy (anonymous):

ooh so otherwise would it be right?

OpenStudy (pulsified333):

I think but it would be 22/36= 11/18

OpenStudy (anonymous):

ok ok :)

OpenStudy (pulsified333):

it says not correct

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

isn't at least 8 = 15/36?

OpenStudy (pulsified333):

how?

OpenStudy (pulsified333):

8=5/36, 9=6/36, 10= 3/36, 11=2/36, 12=1/36

OpenStudy (anonymous):

9 has 4 combinations

OpenStudy (pulsified333):

your right

OpenStudy (pulsified333):

26/36

OpenStudy (anonymous):

15/36 + 11/36 - 6/36 = 20/36

OpenStudy (pulsified333):

5/9?

OpenStudy (anonymous):

wait wait...

OpenStudy (pulsified333):

ok

OpenStudy (anonymous):

yes I think 5/9.

OpenStudy (pulsified333):

I don't get this it says its wrong

OpenStudy (anonymous):

:o is this an online thing?

OpenStudy (pulsified333):

yeah

OpenStudy (anonymous):

ok try 26/36 then

OpenStudy (pulsified333):

i only have 2 more attempts

OpenStudy (anonymous):

:( I don't get it the events are not mutually excessive so it should be 20/36 but if the test writer is dumb it might be 26/36...

OpenStudy (anonymous):

I'm sorry I'm dumb too probably.